OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x+c)2=a 25x-100x-275=0

OpenStudy (anonymous):

i think u miss x in either 25(which should be 25x^2 ) or in the 100(which should be 100x^2)?

OpenStudy (anonymous):

it is 25x^2

OpenStudy (zzr0ck3r):

25x^2-100x-275 25(x^2-4x)-275 25(x-2)^2-275-25(4) 25(x-2)^2-375

OpenStudy (zzr0ck3r):

let me know if you need help understanding what I did.

OpenStudy (anonymous):

shouldnt 25 be to the second power

OpenStudy (anonymous):

yes it is i messed that up

OpenStudy (zzr0ck3r):

@Gunsy do you understand what I did?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

25x2 -100x -275 =0 divide all by 25 x^2 -4x - 11=0 x^2 -4x = 11 dive -4 by 2 which is equal to -2 then square it, it will be 4, now add this number to both sides x^2 -4x +4 = 11+4 (x - 2) ^ 2 = 15 hope it is clear

OpenStudy (anonymous):

that i understood Sam

OpenStudy (anonymous):

i have another one as well 5x^2-100x+455=0 Using the technique of completeing the square to transform the quadratic equation into form (x+c)^2=a

OpenStudy (zzr0ck3r):

if you understood the last one, just do the same thing. If not ask a question. we should not just do them all for you..not that you are asking us to.

OpenStudy (anonymous):

it would be (x-10)^2=9 correct??